OpenStudy (anonymous):

Jennifer makes large cookies. These cookies are 14 inches in diameter. How much icing will she need to cover the top of each cookie? Answer 14π in2 49π in2 7π in2 196π in2

OpenStudy (anonymous):

okay so the radius is half the diameter which is 7. then we have the formula for circumference which is:

OpenStudy (anonymous):

2pir..(or just pi*d) which is 14pi:)

OpenStudy (anonymous):

thank you <3 <3

Directrix (directrix):

----> Answer: 49π in2 Icing on cookies indicates that this is an area of a circle problem. A = pi*r^2 A = pi * (7)^2 A = pi * 49 A = 49 pi square inches